# Effectiveness of the syndrome extraction circuit with flag qubits on IBM quantum hardware

### v1

@@ Abstract
Previous studies have focused on evaluating the quantum error correcting code that does not have flag qubits in real devices such as the Sycamore device from Google, where physical qubits can have more connectivity when compared to the heavy-hexagon structuregoogle_qec1 ; google_qec2 ; neutral ; hh-code ; magic ; d2_qec ; d3_bacon ; d3_ion ; d3_surf . In the case where flag qubits exist, quantum error correction in the current IBM machine has not been explored in terms of increasing the size of the structurerepetition_ibm ; d3_dia ; d3_surper ; ibm_qec ; melbourne . However, in this study, we demonstrate that a syndrome extraction circuit with flag qubits in the repetition code can be successfully implemented on the IBM quantum machine.
@@
A repetition code of distance $d$ consists of a one-dimensional array of data qubits of $n_{data}=d$ and syndrome qubits of $n_{synd}=d-1$. To consider a syndrome extraction circuit with flag qubits in the repetition code, a flag qubit needs to be added between the data and syndrome qubits to ensure they are not adjacent. If a single-flag qubit exists between them, the total number of flag qubits becomes $n_{flag}=2n_{synd}$. However, when two flag qubits are used, the structure requires $n_{flag}=4n_{synd}$ flag qubits. The repetition code can be denoted by $[[n,k,d]]_{f}$, where $n$ and $k$ are the numbers of data qubits and logical qubits, respectively, and $d$ is the distance of the code. Notably, $f=0$ implies no flag qubit; meanwhile, $f=1$ and $f=2$ denote single-flag qubit and double-flag qubits, respectively, between the data and syndrome qubits.
@@
The error in the data qubit that anti-commutes with the stabilizer can be noticed by performing a subroutine quantum circuit, that is, a syndrome extraction circuit. Therefore, syndrome extraction circuits are used to detect the errors. Fig. 2 depicts the Z syndrome extraction circuits in the cases of no-flag qubit ($[[3,1,3]]_{f=0}$), a single-flag qubit ($[[3,1,3]]_{f=1}$), and double-flag qubits ($[[3,1,3]]_{f=2}$) between a data qubit and syndrome qubit. The X syndrome extraction circuits are listed in the supplementary material.
@@
Figure 3: Process example for syndrome extraction round in $[[3,1,3]]_{f=1}$. The figure explains the process of the quantum circuit that samples the result and calculates the syndrome from the measurement result. The quantum circuit consists of three parts: initializing data qubits, R rounds of the syndrome extraction circuit with reset gates on flag and syndrome qubits, and measuring data qubits. The outcomes from each stabilizer extraction circuit are used to obtain the syndrome by employing XOR gates across both temporal and spatial dimensions.
@@
Figure 8: The selected 49 physical qubits for $[[9,1,9]]_{f=2}$ and the error rate of ibm_kyoto. (a) The selected 49 physical qubits for $[[9,1,9]]_{f=2}$ are shown with color dots with respect to their types: data, syndrome, and flag qubits. The blue dots correspond to data qubits, black and red dots represent syndrome, and flag qubits, respectively. (b) The graph shows ECDF of gate errors for selected physical qubits at $[[9,1,9]]_{f=2}$. In this figure, error rates of quantum gates are considered when the initial logical state is $|{0}\rangle_{L}$. The cumulative distribution of each error of ECR, SX($\sqrt{X}$), and readout are plotted as a function of the physical error rate. The dot lines correspond to their mean values.
@@
To implement a syndrome extraction circuit with flag qubits in the repetition code on ibm_kyoto, we select a combination of one-dimensional qubits from a heavy-hexagon structure of 127 qubits. Fig. 7 depicts the error rates of the SX gate, ECR gate, and measurement in ibm_kyotord . The average of the lifetime (T1) and coherence time (T2) are 219.08 $\mu$s and 125.06 $\mu$s, respectively. When a flag qubit is introduced between the data and syndrome qubits, we need 33 qubits in the case of $[[9,1,9]]_{f=1}$, because it uses 16 flag qubits between the data and syndrome qubits. If double-flag qubits exist between the data and syndrome qubits, we need 49 qubits in the case of $[[9,1,9]]_{f=2}$, as depicted in Fig. 8 (a). Fig. 8 (b) shows that the ECR pulse, the measurement, and single qubit gate error rates in the selected qubits for the $[[9,1,9]]_{f=2}$ are expressed in terms of the empirical cumulative distribution function (ECDF).
Fig. 9 depicts the syndrome graph of $[[9,1,9]]_{f=2}$ obtained using the Stim codestim . The weight for each edge in the syndrome graph can be determined by considering the probability of the corresponding error. We assign a lower weight when the probability is high, increasing the likelihood that the MWPM algorithm selects that error as a correction operator. The meaning of color is as follows: As the value of the weight increases, the intensity of the red color increases. Further, as the weight decreases, the intensity of the blue color increases. This means the probability of selecting errors corresponding to edges indicated in blue increases.
@@
Structure $[[n,k,d]]_{f}$
Error Type
Avg. weight
$[[9,1,9]]_{f=0}$
S error
$5.560$
T error
$4.387$
ST error
$5.803$
$[[9,1,9]]_{f=1}$
S error
$5.260$
T error
$2.893$
ST error
$5.487$
$[[9,1,9]]_{f=2}$
S error
$4.874$
T error
$2.038$
ST error
$5.174$
@@
Figure 10: Correlation matrices of repetition code at a distance of nine with different numbers of flag qubits. Space-time correlation matrices for (a) $[[9,1,9]]_{f=0}$, (b) $[[9,1,9]]_{f=1}$, and (c) $[[9,1,9]]_{f=2}$ are shown. The matrices are obtained when the initial logical state is $|1\rangle_{L}$, and the number of rounds in the syndrome extraction is 10. Each pixel can be mapped on the edge of the corresponding syndrome graph. The S, T, and ST errors are visible along the diagonal line. The dashed diagonal lines correspond to multiple T errors occurring consecutively in a row. The axes represent the labeled syndrome node in the syndrome graph in the space-time method. Each syndrome round is divided as $t$, which varies from 0 to 10.
Table 1 lists the average values of the weights for the cases of $[[9,1,9]]_{f=0}$, $[[9,1,9]]_{f=1}$, and $[[9,1,9]]_{f=2}$ with respect to the type of errors: S, T, and ST errors. As listed in Table 1, the weights corresponding to the measurement errors have smaller values regardless of the existence of the flag qubit. This tendency appears in the case of the average error rate of gates of ibm_kyoto. Furthermore, the weights with flag qubits have lower values than those without flag qubits. This implies that the more flag qubits we use, the more likely the errors occur.
@@
Figure 11: Ratio between the number of corresponding sample and total number of sampled data in terms of the number of defect syndrome for (a) $[[9,1,9]]_{f=0}$, (b) $[[9,1,9]]_{f=1}$, and (c) $[[9,1,9]]_{f=2}$. The data is obtained with 10 rounds of syndrome extraction. The probability is calculated by considering Z basis states $|{0}\rangle_{L}$ and $|{1}\rangle_{L}$, meaning the total number of samples is $10^{5}$. The number of defect syndromes is plotted when its probability is more significant than $10^{-5}$ for all distances.
@@
Fig. 10 (a), (b), and (c) depict the correlation matrix of syndrome samples of $[[9,1,9]]_{f=0}$, $[[9,1,9]]_{f=1}$, and $[[9,1,9]]_{f=2}$, respectively, which are displayed in terms of the forms of the space-time method. These figures exhibit the explicit correlation between syndrome nodes in space and time. The probability of detecting two defect syndrome nodes is higher in the presence of flag qubits than in their absence. As we add more flag qubits, we can see a higher probability of an error, which can be broken down into a combination of categorized errors (S, T, and ST error). This implies that the cross-talk effect becomes larger when flag qubits are included. T errors, caused by measurement errors on either syndrome or flag qubits, can occur consecutively over time because they are more frequent than S or ST errors. The phenomenon is observed in all cases, with or without flag qubits, as depicted in Fig. 10 (a), (b), and (c) with the dotted diagonal lines.
Figure 12: Logical error rate of repetition codes. The figure displays the logical error rates of $[[d,1,d]]_{f=0}$, $[[d,1,d]]_{f=1}$, and $[[d,1,d]]_{f=2}$ in terms of syndrome extraction rounds when (a, b, c) hardware-based or (d, e, f) sample-based syndrome graph is used. The logical error rate is shown as the function of the number of the data qubits ($d=3,5,7,9$) with different structures as follows: (a, d) The logical error rate of repetition codes with no flag qubit, (b, e) a single-flag qubit, and (c, f) double-flag qubits.
Structure $[[n,k,d]]_{f}$
Error Type
Avg. weight
$[[9,1,9]]_{f=0}$
S error
$4.416$
T error
$4.350$
ST error
$5.533$
$[[9,1,9]]_{f=1}$
S error
$3.868$
T error
$2.680$
ST error
$5.101$
$[[9,1,9]]_{f=2}$
S error
$2.800$
T error
$1.120$
ST error
$4.321$
@@
Table 2 lists the average weight of the syndrome graph obtained from the correlation matrix for the cases $[[9,1,9]]_{f=0}$, $[[9,1,9]]_{f=1}$, and $[[9,1,9]]_{f=2}$. When the weight is derived from the sampled data, time errors are also more likely to occur, similar to the case with the hardware-based syndrome graph.
Fig. 11 (a), (b), and (c) depict the probability of the samples from the data concerning the number of defect syndromes according to the distance in the cases of no flag qubits, one flag qubit, and two flag qubits, respectively. The probabilities are evaluated by considering two logical qubit states, $|{0}\rangle_{L}$ and $|{1}\rangle_{L}$. The case in which the number of defect syndromes is zero implies that an error detected by the syndrome does not exist. In the case of $[[3,1,3]]_{f=0}$, the number of defect syndromes to be zero is $62.325\%$ in the data, which implies that more than half become error-free. As a flag qubit is added, its probability of being zero decreases ($23.904\%$), and the value is the lowest when double-flag qubits are added ($0.322\%$) for distance three, which is the case of $[[3,1,3]]_{f=2}$. Further, the probability of the samples with an even number of defect syndromes is more significant than that for the number of cases to be odd, regardless of distance and the number of flag qubits. This is because only the S error located at the boundary can produce an odd number of defect syndromes. Notably, the probability of many defect syndromes in the presence of flag qubits is more significant than that without any flag qubits.
@@ Acknowledgements
T1($\mu$s)
T2($\mu$s)
Readout length(ns)
217.69
140.21
1400
1Q gate time(ns)
2Q gate time(ns)
60
660
Readout error
Idling error
$1.90\times 10^{-2}$
$2.79\times 10^{-4}$
1Q gate error
2Q gate error
$2.79\times 10^{-4}$
$8.16\times 10^{-3}$
Table 3: Time and error information for the physical qubits in $[[9,1,9]]_{f=2}$. The table lists the average times of T1, T2, readout, single-qubit gates, and two-qubit gates. The average error rates of readout, idling, single-qubit gates, and two-qubit gates are also displayed.
Here, we describe the hardware specifications of ibm_kyoto. The hardware has a heavy-hexagon structure consisting of transmons of a fixed frequency. ibm_kyoto performs periodic calibrations in qubits and gates. Our study is based on the hardware specification at 2023-12-25-05:39:45 $\sim$ 11:42:34 +00:00. Values can vary over time, even within the period. Table 3 displays a summary of the hardware specifications at the time when we run and obtain data for $|{-}\rangle_{L}$ in $[[9,1,9]]_{f=2}$.
@@
Figure 16: Selected physical qubits for different distances $d=3,5,7,9$. The physical qubits that used in (a) $[[3,1,3]]_{f=2}$, (b) $[[5,1,5]]_{f=2}$, (c) $[[7,1,7]]_{f=2}$, and (d) $[[9,1,9]]_{f=2}$ are shown. The blue, red, and black dots correspond to data, flag, and syndrome qubits.
Figure 17: Gate error rates of the physical qubits in the repetition codes with double-flag qubits. The figure shows the case of (a) $[[3,1,3]]_{f=2}$, (b) $[[5,1,5]]_{f=2}$, (c) $[[7,1,7]]_{f=2}$, and (d) $[[9,1,9]]_{f=2}$. The ECDF is plotted as the function of the error rates of the basis gate set of ibm_kyoto for each case.
@@
We select the physical qubits for the double-flag qubit structure, as depicted in Fig. 16. When the logical qubit is initialized as $|-\rangle_{L}$, the average error rate for the error type for each case is as follows. The average readout error of $[[3,1,3]]_{f=0}$($[9,1,9]]_{f=0}$) is $9.18\times 10^{-3}$($1.06\times 10^{-2}$). The average readout error of $[[3,1,3]]_{f=1}$($[[9,1,9]]_{f=1}$) is $1.11\times 10^{-2}$($1.15\times 10^{-2}$). Meanwhile, the average error rate of the ECR gate of $[[3,1,3]]_{f=0}$($[9,1,9]]_{f=0}$) is $4.80\times 10^{-3}$($5.44\times 10^{-3}$). The average error rate of the ECR gate of $[[3,1,3]]_{f=1}$($[9,1,9]]_{f=1}$) is $5.91\times 10^{-3}$($6.61\times 10^{-3}$). Fig. 17 depicts the qubits and error rates of gates when two flag qubits are introduced between a data qubit and a syndrome qubit. ibm_kyoto provides 127 qubits. Therefore, the qubit selection for the syndrome extraction circuit with flag qubits in the repetition code is not unique. We use the function “transpile” from the qiskit package to choose specific physical qubits that are likely to provide the best performance among the possible combinations. In Fig. 18, the average error rates for each basis gate of the selected physical qubits in the hardware are plotted as a function of the distance and number of flag qubits between the data and syndrome qubits. As more flag qubits are added to the structure, the average error rates for all gates increase.
@@
Figure 19: Procedure for extracting the syndrome from the sampled data. The structure is $[[3,1,3]]_{f=1}$ and the number of the syndrome extraction circuit rounds is three. The initial logical qubit state is $|0\rangle_{L}$. Each syndrome extraction circuit produces six outcomes from the flag and syndrome qubits. Their outcomes can be grouped with red and black boxes. The syndrome bits corresponding to the orange dots can be calculated by checking the parity of two consecutive groups in temporal.
Figure 20: Space-Time correlation matrices obtained in the structures with double-flag qubits. Space-Time correlation matrices are shown in the case of (a) $[[3,1,3]]_{f=2}$, (b) $[[5,1,5]]_{f=2}$, (c) $[[7,1,7]]_{f=2}$, and (d) $[[9,1,9]]_{f=2}$. The data is collected from the experiment with $|-\rangle_{L}$ state. Ten rounds of the syndrome extraction circuit are implemented. The number of samples is 50,000.
Here, we explain the method used to obtain the syndromes from the outcomes of syndrome extraction circuits. When R round syndrome extraction is performed, the syndrome can be broadly divided into three parts in terms of time $T$: 1. at $T=0$, 2. from $T=1$ to $T=(R-1)$, and 3. at $T=R$. The method used to treat this syndrome is depicted in Fig. 19. The diagram depicts an example of a method for obtaining a syndrome with $[[3,1,3]]_{f=1}$ and $R=3$. The initial data qubits are prepared with $|0\rangle_{L}=|000\rangle$. The outcomes from the syndrome extraction circuits can be divided into patches based on the syndrome qubits, and each patch is covered with red and black boxes containing three bits from the syndrome and flag qubits in Fig. 19. The calculation of a syndrome bit involves checking the parity of six values between two consecutive patches. The process is as follows: the system counts the number of occurrences of the digit ‘1’ among these six values. If the count is odd, the syndrome bit is set to ‘1’ otherwise, it is set to ‘0’. This method determines the syndrome bit based on the parity of observed values within a specified temporal range. For a syndrome bit located in the temporal boundary, two data qubits that are neighbors of a patch are used to calculate the syndrome bit. For instance, each syndrome bit in the initial syndrome round at T = 0 utilizes five bits comprising three values from the outcomes of the first round of a syndrome extraction circuit and two values from the initial states of its adjacent data qubits. The process is the same for the syndrome bits in the final rounds ($T=3$), except that the measured states of the data qubits are used instead of their initial states. The method for calculating the syndrome bit-string can be applied similarly to a structure involving double-flag qubits. The only variation lies in the specific number of bits, which is five in the case of double-flag qubits between the data and syndrome qubits within each patch.
@@
Figure 21: Time-Space correlation matrices obtained in the structures with double-flag qubits. Time-Space correlation matrices are shown in the case of (a) $[[3,1,3]]_{f=2}$, (b) $[[5,1,5]]_{f=2}$, (c) $[[7,1,7]]_{f=2}$, and (d) $[[9,1,9]]_{f=2}$. The data is collected from the experiment with $|-\rangle_{L}$ state. Ten rounds of the syndrome extraction circuit are implemented. The number of samples is 50,000.
@@
Figure 22: The probability of defect syndromes. The probability of defect syndromes for the $|{-}\rangle_{L}$ state in the $[[9,1,9]]_{f=2}$ structure is shown in the case of (a) real quantum device, ibm_kyoto, and (b) Stim code. The data is collected from the experiment with $|-\rangle_{L}$ state. Here, the number of syndrome rounds is eleven. The number of samples is 50,000.
The defect rate is defined as the ratio of the number of samples with detected syndromes to the total number of samples. Fig. 22 depicts the defect rate(gray line) from each syndrome qubit for $[[9,1,9]]_{f=2}$ versus the syndrome rounds. The red line represents the average of all defect rates for each syndrome qubits. The first (last) round also uses the initial (measured) state of the data qubits to identify the defect syndromes. Data are obtained by initializing the logical qubit as $|{-}\rangle_{L}$ with 50,000 samples. The number of syndrome extraction cycles is 10. Thus, there are 11 rounds of syndromes. We use data sampled from (a) ibm_kyoto and (b) the Stim code, which adopts the noise model of ibm_kyoto. Both cases have similar behavior regarding the probability of defect syndromes being lower in the first and last rounds than in other rounds and maintaining one value between them. However, those values are not the same, which can be attributed to the oversimplification of each qubit’s error channel as a depolarizing and idling error model. This suggests that the system on the device is more vulnerable to errors than we initially model.

### v2

@@ Abstract
Previous studies have focused on evaluating the quantum error correcting code that does not have flag qubits in real devices such as the Sycamore device from Google, where physical qubits can have more connectivity when compared to the heavy-hexagon structuregoogle_qec1 ; google_qec2 ; neutral ; d2_qec ; d3_bacon ; d3_ion ; d3_surf . In the case where flag qubits exist, quantum error correction in the current IBM machine has not been explored in terms of increasing the size of the structurerepetition_ibm ; d3_dia ; d3_surper ; ibm_qec ; melbourne . However, in this study, we demonstrate that a syndrome extraction circuit with flag qubits in the repetition code can be successfully implemented on the IBM quantum machine.
@@
A repetition code of distance $d$ consists of a one-dimensional array of data qubits of $n_{data}=d$ and syndrome qubits of $n_{synd}=d-1$. To consider a syndrome extraction circuit with flag qubits in the repetition code, a flag qubit needs to be added between the data and syndrome qubits to ensure they are not adjacent. If a single-flag qubit exists between them, the total number of flag qubits becomes $n_{flag}=2n_{synd}$. However, when two flag qubits are used, the structure requires $n_{flag}=4n_{synd}$ flag qubits. The repetition code can be denoted by $[n,k,d]_{f}$, where $n$ and $k$ are the numbers of data qubits and logical qubits, respectively, and $d$ is the distance of the code. Notably, $f=0$ implies no flag qubit; meanwhile, $f=1$ and $f=2$ denote single-flag qubit and double-flag qubits, respectively, between the data and syndrome qubits.
@@
The error in the data qubit that anti-commutes with the stabilizer can be noticed by performing a subroutine quantum circuit, that is, a syndrome extraction circuit. Therefore, syndrome extraction circuits are used to detect the errors. Fig. 2 depicts the Z syndrome extraction circuits in the cases of no-flag qubit ($[3,1,3]_{f=0}$), a single-flag qubit ($[3,1,3]_{f=1}$), and double-flag qubits ($[3,1,3]_{f=2}$) between a data qubit and syndrome qubit. The X syndrome extraction circuits are listed in the supplementary material.
@@
Figure 3: Process example for syndrome extraction round in $[3,1,3]_{f=1}$. The figure explains the process of the quantum circuit that samples the result and calculates the syndrome from the measurement result. The quantum circuit consists of three parts: initializing data qubits, R rounds of the syndrome extraction circuit with reset gates on flag and syndrome qubits, and measuring data qubits. The outcomes from each stabilizer extraction circuit are used to obtain the syndrome by employing XOR gates across both temporal and spatial dimensions.
@@
Figure 8: The selected 49 physical qubits for $[9,1,9]_{f=2}$ and the error rate of ibm_kyoto. (a) The selected 49 physical qubits for $[9,1,9]_{f=2}$ are shown with color dots with respect to their types: data, syndrome, and flag qubits. The blue dots correspond to data qubits, black and red dots represent syndrome, and flag qubits, respectively. (b) The graph shows ECDF of gate errors for selected physical qubits at $[9,1,9]_{f=2}$. In this figure, error rates of quantum gates are considered when the initial logical state is $|{0}\rangle_{L}$. The cumulative distribution of each error of ECR, SX($\sqrt{X}$), and readout are plotted as a function of the physical error rate. The dot lines correspond to their mean values.
@@
To implement a syndrome extraction circuit with flag qubits in the repetition code on ibm_kyoto, we select a combination of one-dimensional qubits from a heavy-hexagon structure of 127 qubits. Fig. 7 depicts the error rates of the SX gate, ECR gate, and measurement in ibm_kyotord . The average of the lifetime (T1) and coherence time (T2) are 219.08 $\mu$s and 125.06 $\mu$s, respectively. When a flag qubit is introduced between the data and syndrome qubits, we need 33 qubits in the case of $[9,1,9]_{f=1}$, because it uses 16 flag qubits between the data and syndrome qubits. If double-flag qubits exist between the data and syndrome qubits, we need 49 qubits in the case of $[9,1,9]_{f=2}$, as depicted in Fig. 8 (a). Fig. 8 (b) shows that the ECR pulse, the measurement, and single qubit gate error rates in the selected qubits for the $[9,1,9]_{f=2}$ are expressed in terms of the empirical cumulative distribution function (ECDF).
Fig. 9 depicts the syndrome graph of $[9,1,9]_{f=2}$ obtained using the Stim codestim . The weight for each edge in the syndrome graph can be determined by considering the probability of the corresponding error. We assign a lower weight when the probability is high, increasing the likelihood that the MWPM algorithm selects that error as a correction operator. The meaning of color is as follows: As the value of the weight increases, the intensity of the red color increases. Further, as the weight decreases, the intensity of the blue color increases. This means the probability of selecting errors corresponding to edges indicated in blue increases.
@@
Structure $[n,k,d]_{f}$ Error Type Avg. weight
$[9,1,9]_{f=0}$ S error $5.560$
T error $4.387$
ST error $5.803$
$[9,1,9]_{f=1}$ S error $5.260$
T error $2.893$
ST error $5.487$
$[9,1,9]_{f=2}$ S error $4.874$
T error $2.038$
ST error $5.174$
@@
Figure 10: Correlation matrices of repetition code at a distance of nine with different numbers of flag qubits. Space-time correlation matrices for (a) $[9,1,9]_{f=0}$, (b) $[9,1,9]_{f=1}$, and (c) $[9,1,9]_{f=2}$ are shown. The matrices are obtained when the initial logical state is $|1\rangle_{L}$, and the number of rounds in the syndrome extraction is 10. Each pixel can be mapped on the edge of the corresponding syndrome graph. The S, T, and ST errors are visible along the diagonal line. The dashed diagonal lines correspond to multiple T errors occurring consecutively in a row. The axes represent the labeled syndrome node in the syndrome graph in the space-time method. Each syndrome round is divided as $t$, which varies from 0 to 10.
Table 1 lists the average values of the weights for the cases of $[9,1,9]_{f=0}$, $[9,1,9]_{f=1}$, and $[9,1,9]_{f=2}$ with respect to the type of errors: S, T, and ST errors. As listed in Table 1, the weights corresponding to the measurement errors have smaller values regardless of the existence of the flag qubit. This tendency appears in the case of the average error rate of gates of ibm_kyoto. Furthermore, the weights with flag qubits have lower values than those without flag qubits. This implies that the more flag qubits we use, the more likely the errors occur.
@@
Figure 11: Ratio between the number of corresponding sample and total number of sampled data in terms of the number of defect syndrome for (a) $[9,1,9]_{f=0}$, (b) $[9,1,9]_{f=1}$, and (c) $[9,1,9]_{f=2}$. The data is obtained with 10 rounds of syndrome extraction. The probability is calculated by considering Z basis states $|{0}\rangle_{L}$ and $|{1}\rangle_{L}$, meaning the total number of samples is $10^{5}$. The number of defect syndromes is plotted when its probability is more significant than $10^{-5}$ for all distances.
@@
Fig. 10 (a), (b), and (c) depict the correlation matrix of syndrome samples of $[9,1,9]_{f=0}$, $[9,1,9]_{f=1}$, and $[9,1,9]_{f=2}$, respectively, which are displayed in terms of the forms of the space-time method. These figures exhibit the explicit correlation between syndrome nodes in space and time. The probability of detecting two defect syndrome nodes is higher in the presence of flag qubits than in their absence. As we add more flag qubits, we can see a higher probability of an error, which can be broken down into a combination of categorized errors (S, T, and ST error). This implies that the cross-talk effect becomes larger when flag qubits are included. T errors, caused by measurement errors on either syndrome or flag qubits, can occur consecutively over time because they are more frequent than S or ST errors. The phenomenon is observed in all cases, with or without flag qubits, as depicted in Fig. 10 (a), (b), and (c) with the dotted diagonal lines.
Figure 12: Logical error rate of repetition codes. The figure displays the logical error rates of $[d,1,d]_{f=0}$, $[d,1,d]_{f=1}$, and $[d,1,d]_{f=2}$ in terms of syndrome extraction rounds when (a, b, c) hardware-based or (d, e, f) sample-based syndrome graph is used. The logical error rate is shown as the function of the number of the data qubits ($d=3,5,7,9$) with different structures as follows: (a, d) The logical error rate of repetition codes with no flag qubit, (b, e) a single-flag qubit, and (c, f) double-flag qubits.
Structure $[n,k,d]_{f}$ Error Type Avg. weight
$[9,1,9]_{f=0}$ S error $4.416$
T error $4.350$
ST error $5.533$
$[9,1,9]_{f=1}$ S error $3.868$
T error $2.680$
ST error $5.101$
$[9,1,9]_{f=2}$ S error $2.800$
T error $1.120$
ST error $4.321$
@@
Table 2 lists the average weight of the syndrome graph obtained from the correlation matrix for the cases $[9,1,9]_{f=0}$, $[9,1,9]_{f=1}$, and $[9,1,9]_{f=2}$. When the weight is derived from the sampled data, time errors are also more likely to occur, similar to the case with the hardware-based syndrome graph.
Fig. 11 (a), (b), and (c) depict the probability of the samples from the data concerning the number of defect syndromes according to the distance in the cases of no flag qubits, one flag qubit, and two flag qubits, respectively. The probabilities are evaluated by considering two logical qubit states, $|{0}\rangle_{L}$ and $|{1}\rangle_{L}$. The case in which the number of defect syndromes is zero implies that an error detected by the syndrome does not exist. In the case of $[3,1,3]_{f=0}$, the number of defect syndromes to be zero is $62.325\%$ in the data, which implies that more than half become error-free. As a flag qubit is added, its probability of being zero decreases ($23.904\%$), and the value is the lowest when double-flag qubits are added ($0.322\%$) for distance three, which is the case of $[3,1,3]_{f=2}$. Further, the probability of the samples with an even number of defect syndromes is more significant than that for the number of cases to be odd, regardless of distance and the number of flag qubits. This is because only the S error located at the boundary can produce an odd number of defect syndromes. Notably, the probability of many defect syndromes in the presence of flag qubits is more significant than that without any flag qubits.
@@ Acknowledgements
T1($\mu$s) T2($\mu$s) Readout length(ns)
217.69 140.21 1400
1Q gate time(ns) 2Q gate time(ns)
60 660
Readout error Idling error
$1.90\times 10^{-2}$ $2.79\times 10^{-4}$
1Q gate error 2Q gate error
$2.79\times 10^{-4}$ $8.16\times 10^{-3}$
Table 3: Time and error information for the physical qubits in $[9,1,9]_{f=2}$. The table lists the average times of T1, T2, readout, single-qubit gates, and two-qubit gates. The average error rates of readout, idling, single-qubit gates, and two-qubit gates are also displayed.
Here, we describe the hardware specifications of ibm_kyoto. The hardware has a heavy-hexagon structure consisting of transmons of a fixed frequency. ibm_kyoto performs periodic calibrations in qubits and gates. Our study is based on the hardware specification at 2023-12-25-05:39:45 $\sim$ 11:42:34 +00:00. Values can vary over time, even within the period. Table 3 displays a summary of the hardware specifications at the time when we run and obtain data for $|{-}\rangle_{L}$ in $[9,1,9]_{f=2}$.
@@
Figure 16: Selected physical qubits for different distances $d=3,5,7,9$. The physical qubits that used in (a) $[3,1,3]_{f=2}$, (b) $[5,1,5]_{f=2}$, (c) $[7,1,7]_{f=2}$, and (d) $[9,1,9]_{f=2}$ are shown. The blue, red, and black dots correspond to data, flag, and syndrome qubits.
Figure 17: Gate error rates of the physical qubits in the repetition codes with double-flag qubits. The figure shows the case of (a) $[3,1,3]_{f=2}$, (b) $[5,1,5]_{f=2}$, (c) $[7,1,7]_{f=2}$, and (d) $[9,1,9]_{f=2}$. The ECDF is plotted as the function of the error rates of the basis gate set of ibm_kyoto for each case.
@@
We select the physical qubits for the double-flag qubit structure, as depicted in Fig. 16. When the logical qubit is initialized as $|-\rangle_{L}$, the average error rate for the error type for each case is as follows. The average readout error of $[3,1,3]_{f=0}$($[9,1,9]_{f=0}$) is $9.18\times 10^{-3}$($1.06\times 10^{-2}$). The average readout error of $[3,1,3]_{f=1}$($[9,1,9]_{f=1}$) is $1.11\times 10^{-2}$($1.15\times 10^{-2}$). Meanwhile, the average error rate of the ECR gate of $[3,1,3]_{f=0}$($[9,1,9]_{f=0}$) is $4.80\times 10^{-3}$($5.44\times 10^{-3}$). The average error rate of the ECR gate of $[3,1,3]_{f=1}$($[9,1,9]_{f=1}$) is $5.91\times 10^{-3}$($6.61\times 10^{-3}$). Fig. 17 depicts the qubits and error rates of gates when two flag qubits are introduced between a data qubit and a syndrome qubit. ibm_kyoto provides 127 qubits. Therefore, the qubit selection for the syndrome extraction circuit with flag qubits in the repetition code is not unique. We use the function “transpile” from the qiskit package to choose specific physical qubits that are likely to provide the best performance among the possible combinations. In Fig. 18, the average error rates for each basis gate of the selected physical qubits in the hardware are plotted as a function of the distance and number of flag qubits between the data and syndrome qubits. As more flag qubits are added to the structure, the average error rates for all gates increase.
@@
Figure 19: Procedure for extracting the syndrome from the sampled data. The structure is $[3,1,3]_{f=1}$ and the number of the syndrome extraction circuit rounds is three. The initial logical qubit state is $|0\rangle_{L}$. Each syndrome extraction circuit produces six outcomes from the flag and syndrome qubits. Their outcomes can be grouped with red and black boxes. The syndrome bits corresponding to the orange dots can be calculated by checking the parity of two consecutive groups in temporal.
Figure 20: Space-Time correlation matrices obtained in the structures with double-flag qubits. Space-Time correlation matrices are shown in the case of (a) $[3,1,3]_{f=2}$, (b) $[5,1,5]_{f=2}$, (c) $[7,1,7]_{f=2}$, and (d) $[9,1,9]_{f=2}$. The data is collected from the experiment with $|-\rangle_{L}$ state. Ten rounds of the syndrome extraction circuit are implemented. The number of samples is 50,000.
Here, we explain the method used to obtain the syndromes from the outcomes of syndrome extraction circuits. When R round syndrome extraction is performed, the syndrome can be broadly divided into three parts in terms of time $T$: 1. at $T=0$, 2. from $T=1$ to $T=(R-1)$, and 3. at $T=R$. The method used to treat this syndrome is depicted in Fig. 19. The diagram depicts an example of a method for obtaining a syndrome with $[3,1,3]_{f=1}$ and $R=3$. The initial data qubits are prepared with $|0\rangle_{L}=|000\rangle$. The outcomes from the syndrome extraction circuits can be divided into patches based on the syndrome qubits, and each patch is covered with red and black boxes containing three bits from the syndrome and flag qubits in Fig. 19. The calculation of a syndrome bit involves checking the parity of six values between two consecutive patches. The process is as follows: the system counts the number of occurrences of the digit ‘1’ among these six values. If the count is odd, the syndrome bit is set to ‘1’ otherwise, it is set to ‘0’. This method determines the syndrome bit based on the parity of observed values within a specified temporal range. For a syndrome bit located in the temporal boundary, two data qubits that are neighbors of a patch are used to calculate the syndrome bit. For instance, each syndrome bit in the initial syndrome round at T = 0 utilizes five bits comprising three values from the outcomes of the first round of a syndrome extraction circuit and two values from the initial states of its adjacent data qubits. The process is the same for the syndrome bits in the final rounds ($T=3$), except that the measured states of the data qubits are used instead of their initial states. The method for calculating the syndrome bit-string can be applied similarly to a structure involving double-flag qubits. The only variation lies in the specific number of bits, which is five in the case of double-flag qubits between the data and syndrome qubits within each patch.
@@
Figure 21: Time-Space correlation matrices obtained in the structures with double-flag qubits. Time-Space correlation matrices are shown in the case of (a) $[3,1,3]_{f=2}$, (b) $[5,1,5]_{f=2}$, (c) $[7,1,7]_{f=2}$, and (d) $[9,1,9]_{f=2}$. The data is collected from the experiment with $|-\rangle_{L}$ state. Ten rounds of the syndrome extraction circuit are implemented. The number of samples is 50,000.
@@
Figure 22: The probability of defect syndromes. The probability of defect syndromes for the $|{-}\rangle_{L}$ state in the $[9,1,9]_{f=2}$ structure is shown in the case of (a) real quantum device, ibm_kyoto, and (b) Stim code. The data is collected from the experiment with $|-\rangle_{L}$ state. Here, the number of syndrome rounds is eleven. The number of samples is 50,000.
The defect rate is defined as the ratio of the number of samples with detected syndromes to the total number of samples. Fig. 22 depicts the defect rate(gray line) from each syndrome qubit for $[9,1,9]_{f=2}$ versus the syndrome rounds. The red line represents the average of all defect rates for each syndrome qubits. The first (last) round also uses the initial (measured) state of the data qubits to identify the defect syndromes. Data are obtained by initializing the logical qubit as $|{-}\rangle_{L}$ with 50,000 samples. The number of syndrome extraction cycles is 10. Thus, there are 11 rounds of syndromes. We use data sampled from (a) ibm_kyoto and (b) the Stim code, which adopts the noise model of ibm_kyoto. Both cases have similar behavior regarding the probability of defect syndromes being lower in the first and last rounds than in other rounds and maintaining one value between them. However, those values are not the same, which can be attributed to the oversimplification of each qubit’s error channel as a depolarizing and idling error model. This suggests that the system on the device is more vulnerable to errors than we initially model.